# organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France

### v1

@@ 2 Induced path polytope of chordal graphs
Which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. One can repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{u}-y^{*}(\delta(v)))$ yields the desired conic combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
@@
[…]
Figure 3: Induced paths considered in the proof of Theorem 2.2
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. These constraints have been added to give easy initializations of the induction. The proof of Theorem 2.2 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
@@
A graph $G$ is chordal if and only if its extended space induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets, we characterize which do.
@@
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\in N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K)]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K)]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K,\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 8 concludes:
@@
    $\xi^{u,v}\ \forall uv\in E\setminus\delta(w)$
@@
    $\xi^{u,w,v}\ \forall uw\in\delta(w)\setminus\delta(K)\setminus\delta(C(K))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K)]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to be spanning the vertices. Now, match each edge of $\delta(w,C(K))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K)|=|\delta(w,C(K))|$. Let $uw$ be in $\delta(w,C(K))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non adjacent hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such path will contain two edges of $\delta(w,C(K))$ making it hard to build an indentity of wanted size over these components.
We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and ”?” stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof. The rank of $H$ is unkown, however, we know that the restriction of the last $|\delta(w,C(K))|$ points to elements of $G\setminus\delta(w,C(K))$ contains a incidence matrix of a non bipartite graph $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K)|$.
Linear independence is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the latter set of points forms a submatrix of rank $|C(K)|$ which are the only ones involving elements of $\delta(w,C(K))$. ∎
@@
The characterization of Theorem 2.7 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, if $K\subseteq K^{\prime}$ then we have $C(K^{\prime})\subseteq C(K)$, for this reason, we say that a clique defines an orbit if there exists no $K^{\prime}$ strictly containing $K$ such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
@@
Given a chordal graph $G$, the set of non-dominated inequalities inequalities (5) has polynomial size.
@@
Let $K_{H}$ be a maximal clique of $H$. It is partitioned into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{k}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
@@
The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.7 and 2.8 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.7 implies that its associated inequalities are not facet defining.
@@ Conclusion
In this paper, we propose extended formulations for the induced tree and path polytopes of chordal graphs of compact size, implying the polynomial solvability of finding a maximum vertex and edge weighted such induced subgraph in chordal graphs. Our proofs rely heavily on the structure of the induced trees and paths cones; it may be interesting to investigate the structures of the linear programs whose constraint sets correspond to incidence vectors of induced paths and trees.
Acknowledgment
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support.

### v2

@@ 2.1 Induced paths of chordal graphs and Hilbert basis
which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. Repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired integer combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
As induced trees of $C_{4}$ coincide with its induced path, Theorem 2.8 is not a characterization of the graph for which the induced path extended incidence vectors have IDP.
2.2 Extended formulation for the induced path polytope
In this section, we discuss the consequences of Theorem 2.8 and use it to give a compact description of the induced path polytope of chordal graphs.
Corollary 2.9.
The induced path cone and polytope of chordal graphs both have IDP.
Similarly to what we do for the induced tree polytope, we exhibit a description of the induced path polytope of chordal graphs. We denote by $\mathcal{P}(G)$ the polytope defined as the intersection of $\mathcal{C}_{P}(G)$ and the hyperplane:
$\{(x,y)~|~x(V)-y(E)=1\}.$ (11)
Theorem 2.10.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary.
Proof. $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11). Theorem 2.8 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11), which concludes the proof.
@@ 2.2 Extended formulation for the induced path polytope
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. The proof of Theorem 2.10 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
@@
A graph $G$ is chordal if and only if its induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets; we characterize which ones do.
@@
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\subseteq N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K\cup\{w\})]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 11 concludes. We consider $K_{w}=K\cup\{w\}$. The points are as follows:
@@
    $\xi^{\{u,v\}}\ \forall uv\in E\setminus\delta(w)$
@@
    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus(\delta(K)\cup\delta(C(K_{w})))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K_{w}))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to span the vertices. Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$. Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such a path will contain two edges of $\delta(w,C(K_{w}))$, making it impossible to continue a diagonal equal to 1. We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and "?" stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
$H$ is similar to the incidence matrix of the non bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
Full row rank is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the last set of points forms a submatrix of rank $|C(K_{w})|$ which are the only ones involving elements of $\delta(w,C(K_{w}))$. This implies that the matrix has rank $|V|+|E|$, therefore the inequality is facet defining. ∎
@@
The characterization of Theorem 2.12 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$. If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$, for this reason, we say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
@@
Given a chordal graph $G$, the set of non-dominated inequalities (5) has polynomial size.
@@
Let $K_{H}$ be a maximal clique of $H$. It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{K}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$. Summing over all $w\in V$, the total number of non-dominated inequalities (5) is $O(|V|^{2})$.
@@
There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$. The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.12 and 2.13 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.12 implies that its associated inequalities are not facet defining.
@@ Conclusion
In this paper, we show that the induced tree and path extended incidence vectors of chordal graphs form a Hilbert basis. We use this fact to give a compact linear system of inequalities whose binaricity characterizes chordality. It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs. In general, showing that a set of vectors $\mathcal{V}$ forms a Hilbert basis is used to show that a linear system whose rows are elements of $\mathcal{V}$ is TDI. In this paper, we are not considering the programs whose rows are extended induced path or induced tree of chordal graphs. Finding some problems that are naturally formulated as integer linear programs whose linear constraint set has such a structure would lead to TDI-ness results.
Statements and Declarations
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support and declares that he has no conflicts of interest.